### v2

# Teaching Code LLMs to Use Autocompletion Tools in Repository-Level Code Generation

Abstract
Recent code large language models (LLMs) have shown promising performance in generating standalone functions. However, they face limitations in repository-level code generation due to their lack of awareness of repository-level dependencies (e.g., user-defined attributes), resulting in dependency errors such as undefined-variable and no-member errors. In this work, we introduce ToolGen, an approach that integrates autocompletion tools into the code LLM generation process to address these dependencies. ToolGen comprises two main phases: Trigger Insertion and Model Fine-tuning (Offline), and Tool-integrated Code Generation (Online). During the offline phase, ToolGen augments functions within a given code corpus with a special mark token, indicating positions to trigger autocompletion tools. These augmented functions, along with their corresponding docstrings, are then used to fine-tune a selected code LLM. In the online phase, ToolGen iteratively generates functions by predicting tokens step-by-step using the fine-tuned LLM. Whenever a mark token is encountered, ToolGen invokes the autocompletion tool to suggest code completions and selects the most appropriate one through constrained greedy search.
We conduct comprehensive experiments to evaluate ToolGen’s effectiveness in repository-level code generation. To facilitate this evaluation, we create a benchmark comprising 680 real-world code repositories and introduce two new repository-level metrics: Dependency Coverage and Static Validity Rate. The results demonstrate that ToolGen significantly improves Dependency Coverage by 15.2% to 45.8% and Static Validity Rate by 10.9% to 42.2% across three distinct code LLMs, while maintaining competitive performance in widely-recognized similarity metrics such as BLEU-4, CodeBLEU, and Edit Similarity. Furthermore, our generalizability evaluation confirms ToolGen’s consistent performance when applied to diverse code LLMs, including various model architectures and scales.
I Introduction
Code generation has been a longstanding focal point in the field of software engineering. Recent advancements have introduced a variety of code large language models (LLMs) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] constructed upon the Transformer model architecture [14]. These models are either pre-trained or fine-tuned on extensive code corpora, enabling them to automatically generate code based on provided natural language descriptions. These code LLMs have demonstrated notable effectiveness in the generation of code blocks or functions. For instance, CodeLlama [12], built upon the foundational Llama2 model [15], has achieved state-of-the-art results among open code LLMs (e.g., CodeGen [8] and StarCoder [11]), on benchmarks like HumanEval [3] and MBPP [16] that focus on standalone functions.
However, it is crucial to emphasize that in real-world code repositories, more than 70% of functions are not standalone [17]. Code LLMs encounter significant challenges when generating such real-world functions, primarily because they cannot be aware of repository-level dependencies, such as user-defined functions and attributes, during the code generation process [17]. This limitation often leads to the generation of code with dependency errors, including undefined-variable and no-member errors. These errors impede the usability and effectiveness of the code LLMs [18]. For example, consider the scenario depicted in Figure 1. A code LLM (e.g., CodeLlama) might incorrectly predict “_updates” after generating “... self.”, resulting in a no-member error because the object “self” does not possess an attribute named “_updates”.
Meanwhile, modern Integrated Development Environments (IDEs) take a different approach, which typically incorporates code autocompletion tools based on program analysis. These tools, like Jedi [19], leverage their ability to analyze the current incomplete function’s state and project context to provide valid completion recommendations. This includes suggestions for accessible variables, attributes, and functions. For instance, when encountering “self.” in Figure 1, Jedi can infer and recommend 68 accessible attributes defined within “self”, including the target suggestion “_registered_updates”. Therefore, if we can seamlessly switch between code LLMs and the use of autocompletion tools, we have the potential to significantly reduce the occurrence of dependency errors in repository-level code generation.
Figure 1: Illustrative Example of LLM Prediction and Tool Completion
In fact, recent research has delved into the integration of external tools into the generation process of LLMs to mitigate their limitations in constrained generation scenarios. One noteworthy example is ToolFormer [20], which creates an augmented dataset to instruct LLMs on invoking existing arithmetic calculators. This integration effectively reduces errors in generated text involving arithmetic calculations. Building upon ToolFormer’s inspiration, Zhang et al. [21] introduce ToolCoder, an approach designed to teach LLMs how to utilize information-retrieval-based (IR-based) API search tools during the code generation process. While ToolCoder targets the generation of functionally correct standalone functions and demonstrates promising results, the integrated IR-based API search tools do not consider repository-level dependencies, limiting their potential in resolving dependency errors. Another relevant example of harnessing external tools is Repilot [22], which leverages code completion tools to filter out impractical suggestions made by LLMs in the context of automatic program repairing (APR). Unlike repository-level code generation, Repilot’s primary focus is on generating valid bug-fix patches rather than entire functions.
In this work, we aim at integrating program-analysis-based code autocompletion tools into the generation process of code LLMs. Achieving the incorporation presents two key challenges. (i) Determining when to trigger the invocation of autocompletion tools during the generation process: The generation process of LLMs is a step-by-step decoding process where each subsequent token is predicted based on previous tokens. In general, a function consists of dozens or even hundreds of tokens, making it impractical to invoke code autocompletion tools at every decoding step. In the case of tools like ToolFormer and ToolCoder, ChatGPT is employed to augment the training corpus by introducing special tokens into the text or code to mark positions where tool invocation is needed. After training on this augmented corpus, LLMs can predict the special token at the appropriate step, thereby triggering tool invocation. However, this ChatGPT-based augmentation method is less effective for repository-level code generation due to the presence of repository-level dependencies. The special token must be precisely inserted at positions involving such dependencies, such as when accessing user-defined variables. (ii) Selecting the target suggestion from the recommended completions of autocompletion tools: Different from tools like arithmetic calculation or API search integrated into ToolFormer and ToolCoder, which return a single result for each invocation, autocompletion tools often provide multiple completion suggestions sorted alphabetically. For instance, as depicted in Figure 1, Jedi returns a list of 68 completion suggestions (excluding builtin attributes), with the target suggestion being the 46th one in the list. Consequently, after invoking autocompletion tools, it is essential to assess the suggestions based on the generated code and select the most appropriate one. Furthermore, this selection process needs to be seamlessly integrated into the code generation process to ensure efficiency and coherence.
To tackle the challenges, we propose ToolGen, an approach to integrate autocompletion tools into the generation process of code LLMs to support repository-level code generation. ToolGen has two main phases: Trigger Insertion and Model Fine-tuning (Offline), and Tool-integrated Code Generation (Online). In the offline phase, ToolGen analyzes source files within a corpus of code repositories, creating abstract syntax trees (ASTs) and extracting function definitions. It augments these functions by inserting a special token, <COMP>, signifying the positions to trigger autocompletion tools. The insertion positions are established by navigating through the functions and identifying the identifiers that can be recommended by autocomplete tools. These augmented functions, paired with their respective docstrings, are then employed to fine-tune a selected code LLM. In the online phase, ToolGen iteratively constructs a function based on a provided docstring by predicting tokens step-by-step through the fine-tuned LLM. Whenever a <COMP> token is encountered, ToolGen invokes the autocompletion tool to suggest code completions, drawing from the current repository context. Subsequently, it identifies the most appropriate suggestion through a constraint greedy search algorithm, appending this selected suggestion to the current tokens. This process continues as it predicts tokens until a specified termination condition is satisfied.
We conduct extensive experiments to evaluate the effectiveness of ToolGen in repository-level code generation. To facilitate this evaluation, we first construct a benchmark, which includes 12,744 Python functions from 680 real-world code repositories and defines two new repository-level metrics, namely Dependency Coverage and Static Validity Rate. Dependency Coverage quantifies the proportion of repository-level dependencies present in ground-truth functions and successfully covered by the generated functions, while Static Validity Rate measures the percentage of generated functions that pass a dependency error check. The evaluation results demonstrate that ToolGen exhibits comparable performance in widely-recognized similarity metrics such as BLEU-4, CodeBLEU, and Edit Similarity. Importantly, ToolGen achieves significant improvements in Dependency Coverage, ranging from 15.2% to 45.8%, and Static Validity Rate, spanning from 10.9% to 42.2%, across three distinct code LLMs. Moreover, the results from our generalizability evaluation confirm that ToolGen consistently performs well across a variety of code LLMs, with different model architectures and scales.
In summary, this paper presents the following key contributions:
  • ToolGen, an novel approach that seamlessly integrates autocompletion tools into the generation process of code LLMs, thereby enhancing repository-level code generation.
  • An Augmented Dataset, which comprises 256,870 Python functions sourced from a diverse selection of 12,231 code repositories. Each function is augmented with a special token, <COMP>, which signifies positions suitable for invoking autocompletion tools.
  • An Evaluation Benchmark, which encompasses 12,744 Python functions drawn from 680 real-world code repositories, along with the introduction of two novel repository-level metrics: Dependency Coverage and Static Validity Rate.
  • Extensive Experimental Results, which affirm the efficacy of ToolGen in repository-level code generation. ToolGen demonstrates substantial improvements in Dependency Coverage, ranging from 15.2% to 45.8%, and Static Validity Rate, spanning from 10.9% to 42.2%, across three distinct code LLMs.
II Preliminaries
II-A Code LLMs
Typically, there are two main categories of code LLMs that can be employed for code generation. These categories include decoder-only models and encoder-decoder models, each of which conducts the code generation process base on a given docstring as outlined below:
  • Decoder-only Models: Illustrated in Figure 1(a), decoder-only code LLMs, such as CodeGPT [23] and CodeLlama [12], consist solely of a decoder component derived from the Transformer architecture [14]. An employed decoder-only model first tokenizes the input docstring into a sequence of tokens. Subsequently, it feeds this token sequence into the model’s decoder and proceeds to predict a function token-by-token, based on the context provided by the docstring and previously predicted tokens.
  • Encoder-Decoder Models: As depicted in Figure 1(b), encoder-decoder code LLMs, such as CodeT5 [4] and CodeT5+ [5], encompass both the encoder and decoder components of the Transformer architecture. In this case, the employed model also tokenizes the docstring into a token sequence, but the sequence is first processed by the model’s encoder. The model’s decoder is then tasked with predicting a function token-by-token, relying on the representation produced by the encoder and the context provided by the preceding tokens.
(a) Decoder-only Model
(b) Encoder-Decoder Model
Figure 2: Decoder-only Model and Encoder-Decoder Model
On top of the standard generation process, to ensure that the employed code LLM can recognize and predict the special token <COMP>, we initially incorporate this token into the LLM’s vocabulary, denoted as $\mathbb{V}_{llm}$. Formally, this addition results in an expanded vocabulary represented as:
[…]
For the employed code LLM, within the generation process, we define its tokenization process as a procedure:
[…]
Here, $\Sigma_{char}^{*}$ represents a character sequence of either a docstring or a code snippet, and $\mathbb{V}^{*}$ corresponds to the resulting sequence of tokens drawn from $\mathbb{V}$.
The next token prediction involved in each step is defined as a procedure:
[…]
In this context, the two input token sequences ($\mathbb{V}^{*}$) represent a docstring and an incomplete function, respectively, while $[0,1]^{|\mathbb{V}|}$ signifies a probability distribution encompassing $|\mathbb{V}|$ probabilities $[0,1]$. Here, $|\mathbb{V}|$ is the size (token numbers) of the vocabulary $\mathbb{V}$.
Example: In Figure 1, CodeLlama takes the description “Register updates…” and the incomplete function “… self.” as inputs. It then performs a prediction, generating a probability distribution of size $|\mathbb{V}|$, wherein the token “_updates” exhibits the highest probability among all tokens.
II-B Autocompletion Tools
An autocompletion tool takes a code repository and a caret position (defined as a tuple containing source file, line number, and column number) as input and provides a list of completion suggestions. We define this completion process as a procedure:
[…]
Here, $\Sigma_{repo}$ and $\Sigma_{pos}$ respectively represent the domains of code repositories and caret positions, $\Sigma_{iden}$ encompasses all possible identifiers such that $\Sigma_{iden}^{*}$ is a list of identifiers. It’s worth noting that autocompletion tools often provide a wide range of completion suggestions, including keywords and partial identifiers. In our context, we focus solely on identifier-level completions, as keywords are relatively straightforward for code LLMs to predict, and partial identifiers are encompassed by identifier-level completions.
Example: In Figure 1, when provided with the code repository and caret position, Jedi is capable of generating 86 completion suggestions for the incomplete function "… self.".
III Approach
In this section, we elaborate on our approach named ToolGen to integrate autocompletion tools into the generation process of code LLMs to support repository-level code generation.
III-A Overview
Figure 3 presents an overview of ToolGen, which consists of two main phases, namely (i) Trigger Insertion and Model Fine-tuning (Offline) and (ii) Tool-integrated Code Generation (Online).
Figure 3: Approach Overview of ToolGen
In trigger insertion and model fine-tuning, ToolGen parses each source file in the given code repositories into an abstract syntax tree (AST) and then extract extracts function definitions from the AST; For each extract function definition, ToolGen then utilizes an autocompletion tool to augment it with the special token <COMP> to mark the positions to invoke the tool, and then assembles a pair of docstring and augmented function; After process all code repositories, ToolGen employs the resulting pairs of docstrings and augmented functions to fine-tune a code LLM, resulting in a fine-tuned code LLM that can predict <COMP> at suitable positions to trigger the autocompletion tool.
In tool-integrated code generation, ToolGen generates a token sequence to form a function by an iterative process, in which, at each step, one or multiple tokens are yielded by the fine-tuned code LLM and the employed autocompletion tool. At certain step, ❶ the fine-tuned code LLM takes the given docstring and the incomplete function as inputs and predicts the next token; The predicted token is appended to the incomplete function; ❷ If the predicted token equals <COMP>, the autocomplete tool is triggered and a list of completion suggestions is returned based on the current repository context; ❸ ToolGen then selects the most suitable one from the suggestions with the fine-tuned code LLM and appends the selected suggestion to the incomplete function.
III-B Trigger Insertion and Model Fine-tuning
III-B1 Trigger Insertion
We employ a trigger insertion method to facilitate the learning process of code LLMs in determining when to utilize autocompletion tools during code generation. In this method, the special token <COMP> is inserted at specific locations within code functions, indicating when autocompletion tools should be triggered.
Given a code repository $\mathcal{R}$, we traverse each source file $file$ within it based on the file’s suffix (e.g., .py for Python) and then proceed to analyze the functions defined in the source file. To achieve this, we parse the source file into an abstract syntax tree (AST), where the functions are represented as function-definition nodes. Each function-definition node contains multiple AST-tokens, which are smallest individual units, such as keywords, identifiers, literals, operators, and punctuators, within programming language syntax. Note that these AST-tokens differ from the tokens in the LLM’s vocabulary $\mathbb{V}_{llm}$. Typically, an AST-token comprises one or more tokens from $\mathbb{V}_{llm}$. For example, the AST-token “_registered_updates” consists of six tokens in vocabulary of CodeLlama, i.e., [“_”, “register”, “ed”, “_”, “up”, “dates”].
For each function within the source file, we identify its corresponding function-definition node, denoted as $node$, and apply Algorithm 1 to it. The purpose of this algorithm is to traverse the function body and identify specific identifiers that are eligible for suggestions by autocomplete tools. Subsequently, the special token <COMP> is inserted in front of these chosen identifiers. More specifically, as the algorithm iterates through each AST-token $t$ within the function body $node.body$ (line 2), it performs two crucial checks. First, it employs the isIdentifier procedure to determine whether $t$ is an identifier. Second, it verifies that $t$ is not a built-in attribute, such as “__dict__” in Python, using the isBuiltin procedure. These conditions are essential because dependency errors often arise from user-defined attributes categorized as identifiers rather than other AST-tokens like language keywords. Additionally, these checks prevent the insertion of <COMP> at positions where the code LLM can confidently predict the following tokens, thus minimizing unnecessary tool invocations. When both conditions are met, the algorithm updates the caret position $\mathcal{P}$ to the start position of $t$ (line 4) and invokes the autocompletion tool to obtain a list of completion suggestions, denoted as $\mathbb{C}$ (line 5). If $\mathbb{C}$ contains $t$, indicating that the tool can propose the desired identifier, the special token <COMP> is inserted before $t$ to mark the position for triggering the autocompletion tool (lines 6-7). Upon executing the algorithm, we obtain the augmented function code $\mathcal{F}_{aug}$.
Input: Repository $\mathcal{R}$, Source File $file$, Function-definition node $node$
Output: Augmented function $\mathcal{F}_{aug}$
$\mathcal{F}_{aug}\leftarrow\textsc{getSignature}(node)$   // signature
1 for $t$ in $node.body$ do
2       if $\textsc{isIdentifier}(t)$ and not $\textsc{isBuiltin}(t)$ then
3             $\mathcal{P}=(file,t.start\_line,t.start\_column)$
4             $\mathbb{C}\leftarrow\textsc{tool-complete}(\mathcal{R},\mathcal{P})$
5             if $t\in\mathbb{C}$ then
[…]
7            $\mathcal{F}_{aug}.append(t)$
8
Algorithm 1 Trigger Insertion
Next, we assemble a tuple $(\mathcal{D},\mathcal{F}_{aug})$, in which $\mathcal{D}$ corresponds to the docstring of the parsed function. Note that functions lacking corresponding docstrings are omitted from our process as our repository-level code generation relies on textual descriptions as input. Once we complete the processing of all code repositories, we accumulate an augmented dataset that contains a substantial number of these data tuples.
Example: In Figure 4, we showcase an augmented function that contains four instances of the special token <COMP>. These tokens have been inserted at positions where the desired identifiers, namely “updates”, “_registered_updates”, “add”, and “update”, are found within the suggestion lists of the autocompletion tool.
Figure 4: Augmented Function
III-B2 Model Fine-tuning
During the fine-tuning process, we supply the collected docstrings and augmented functions to optimize the parameters of the employ code LLM (base model), adhering to established practices in code generation tasks. Specifically, for each pair consisting of a docstring $\mathcal{D}$ and an augmented function $\mathcal{F}_{aug}$, both are tokenized into sequences of tokens and subsequently fed into the base model to undergo the token-by-token generation process described in Section II-A. At each step, a cross-entropy loss is computed between the predicted probability distribution of the next token and the ground-truth next token present in $\mathcal{F}_{aug}$.
In the case of code LLMs with an extensive number of parameters, such as CodeLlama-7B with 7 billion parameters, fine-tuning all parameters becomes computationally challenging due to resource limitations. To address this, we employ Low-Rank Adaptation (LoRA) [24] as a parameter-efficient fine-tuning technique. LoRA relies on low-dimensional representations and a freeze-and-inject strategy, where the majority of the model parameters remain fixed, and trainable low-rank matrices are introduced into specific transformer layers, particularly the projection matrices within the attention module, to approximate weight updates.
III-C Tool-integrated Code Generation
Based on the fine-tuned code LLM and the employed autocompletion tool, we perform a tool-integrated code generation process that is aware to the repository-level dependencies.
III-C1 Overall Process
Algorithm 2 outlines the overall tool-integrated generation process, comprising three crucial parts based on the fine-tuned code LLM and the employed autocompletion tool: ❶ Next Token Prediction, ❷ Code Autocompletion, and ❸ Suggestion Selection. This algorithm takes a code repository $\mathcal{R}$, an insertion position $\mathcal{P}$, and a docstring $\mathcal{D}$ as inputs and follows an iterative process to generate a token sequence, ultimately constructing a function denoted as $\mathcal{F}$. Here, the tokens are drawn from the expanded vocabulary $\mathbb{V}$ defined in Equation 1.
[…]
If $tok$ corresponds to the special token <COMP>, the autocompletion tool is triggered to provide a list of completion suggestions denoted as $\mathbb{C}$. These suggestions are produced based on the code repository $\mathcal{R}$ and the caret position $\mathcal{P^{\prime}}$ after inserting $\mathcal{F}$ at $\mathcal{P}$ (lines 9-11). Notably, when $\mathcal{F}$ is inserted using the insert procedure, any <COMP> tokens within it are removed to prevent syntax errors. The fine-tuned code LLM is then employed to assess the completion suggestions and select the most suitable one for $\mathcal{F}$ by the llm-select procedure (line 12). The tokens from the selected suggestion are concatenated to $\mathcal{F}$.
Example: In the case of the incomplete code snippet shown in Figure 1, Algorithm 2 predicts the next token as <COMP> through the fine-tuned code LLM. This prediction triggers the autocompletion tool. Subsequently, the resulting completion suggestions are fed into the llm-select procedure, which determines the most appropriate suggestion.
Input: Repository $\mathcal{R}$, Docstring $\mathcal{D}$, Insertion Position $\mathcal{P}$
Output: Function $\mathcal{F}$
1 $\mathcal{D}\leftarrow\textsc{llm-tokenize}(\mathcal{D})$
[…]
3 while true do
       /* ❶ Next Token Prediction */
4       $\bm{p}^{|\mathbb{V}|}\leftarrow\textsc{llm-predict}(\mathcal{D},\mathcal{F})$
5       $tok\leftarrow\textsc{argmax}(\mathbb{V},\bm{p}^{|\mathbb{V}|})$
6       $\mathcal{F}\leftarrow\mathcal{F}\oplus[tok]$
[…]
8             break
      /* ❷ Code Autocompletion */
[…]
10             $\mathcal{P^{\prime}}\leftarrow\textsc{insert}(\mathcal{P},\mathcal{F})$
11             $\mathbb{C}\leftarrow\textsc{tool-complete}(\mathcal{R},\mathcal{P^{\prime}})$
             /* ❸ Suggestion Selection */
[…]
13
14
Algorithm 2 Tool-integrated Code Generation
III-C2 Completion Suggestion Selection
Algorithm 3 provides a description of the llm-select procedure, which is called within Algorithm 2. To begin, it tokenizes each completion in $\mathbb{C}$ into a sequence of tokens from $\mathbb{V}$ using the code LLM’s tokenizer (via the llm-tokenize procedure) and inserts this token sequence into a prefix tree [26], denoted as $trie$ (lines 1-5). Each node in the tree possesses four properties: $node.token$, $node.tok\_idx$, $node.children$, and $node.is\_terminal$, indicating the token stored in the node, the index of the stored token in $\mathbb{V}$, the child nodes of the current node, and whether the node corresponding to the terminal of a token sequence. The root node, $trie.root$, is a unique node that stores $\epsilon$, signifying an empty string. Every path from $trie.root$ to a terminal node corresponds to a token sequence from $\mathbb{C}$. As an illustration, Figure 5 presents the prefix tree corresponding to the 68 completion suggestions shown in Figure 1. In this example, nodes enclosed in blue boxes indicate the terminals of token sequences.
Figure 5: Example Prefix Tree
1 Procedure llm-select($\mathcal{D},\mathcal{F}$, $\mathbb{C}$):
       $trie\leftarrow Trie()$   // prefix tree
2       for $comp$ in $\mathbb{C}$ do
3             $seq\leftarrow\textsc{llm-tokenize}(comp)$
4             $trie.insert(seq)$
5
6      $node\leftarrow trie.root$
7       while not $node.is\_terminal$ do
8             $\bm{m}^{|\mathbb{V}|}\leftarrow\bm{0}$
9             for $child$ in $node.children$ do
10                   $\bm{m}^{|\mathbb{V}|}[child.tok\_idx]\leftarrow 1$
11            $\bm{p}^{|\mathbb{V}|}\leftarrow\textsc{llm-predict}(\mathcal{D},\mathcal{F})$
12             $\bm{p}^{|\mathbb{V}|}\leftarrow\bm{p}^{|\mathbb{V}|}\odot\bm{m}^{|\mathbb{V}|}$
13             $tok\leftarrow\textsc{argmax}(\mathbb{V},\bm{p}^{|\mathbb{V}|})$
14             $\mathcal{F}\leftarrow\mathcal{F}\oplus tok$
15             for $child$ in $node.children$ do
16                   if $child.token=tok$ then
17                         $node\leftarrow child$
18                         break
19
20
21
Algorithm 3 Suggestion Selection based on Constraint Greedy Search
Subsequently, the algorithm proceeds to select a path in $trie$ in a greedy fashion, based on predictions made by the fine-tuned code LLM, and appends the token sequence associated with the chosen path to the incomplete function $\mathcal{F}$ (lines 6-13). Specifically, the algorithm initiates a node pointer, denoted as $node$, with the root node $trie.root$ (line 6). A loop continues until the pointer $node$ reaches a terminal node (line 7). Within this loop, a $|\mathbb{V}|$-dimensional mask vector, denoted as $\bm{m}^{|\mathbb{V}|}$, is generated based on the children of the current node (lines 8-10). In $\bm{m}^{|\mathbb{V}|}$, only positions corresponding to the $tok\_idx$ property of the children of $node$ are assigned a value of 1, while all other positions are set to 0. Subsequently, the fine-tuned code LLM is employed to predict a probability distribution, $\bm{p}^{|\mathbb{V}|}$ (line 11). This predicted distribution is then element-wise multiplied by the mask vector $\bm{m}^{|\mathbb{V}|}$, effectively setting the probability of tokens not in the children of $node$ to 0. The next token, $tok$, is selected from $\mathbb{V}$ based on the highest probability in $\bm{p}^{|\mathbb{V}|}$ using the argmax function and is appended to the current incomplete function $\mathcal{F}$ (lines 13-14). Finally, the node pointer is updated to point to the child of $node$ whose stored token matches the selected token $tok$ (line 15-18).
Example: For the prefix tree illustrated in Figure 5, the llm-select procedure iteratively selects the next tokens within the tree, guided by the LLM’s predictions. This iterative process results in the inclusion of tokens corresponding to the suggestion “_registered_updates”, which are found along the green path, being appended to the incomplete function.
IV Evaluation Setup
To evalute the effectiveness of ToolGen in repository-level code generation, we conduct a comprehensive set of experiments.
IV-A Research Questions
We formulate the following research questions to guide our evaluation:
  • RQ1 - Similarity Metrics: How closely does the code generated by ToolGen align with the ground truth when assessed using common similarity metrics?
  • RQ2 - Repository-level Metrics:
    • RQ2.a: What extent does the generated code cover repository-level dependencies, including user-defined functions and attributes?
    • RQ2.b: To what degree can ToolGen reduce errors, such as no-member and undefined-variable errors, arising from repository-level dependencies?
  • RQ3 - Generalizability: Is ToolGen effective in code generation when applied to different code LLMs?
IV-B Implementation
Although ToolGen is designed to be language-agnostic, our current focus is on developing a Python-specific prototype of ToolGen.
Base Model. In ToolGen, we explore the utilization of three distinct code LLMs to encompass diverse model architectures and parameter scales. These code LLMs demonstrate impressive performance in code generation and have found extensive utilization in prior studies [23, Compilable, 4, 5, 12] for fine-tuning and evaluation.
  • CodeGPT: CodeGPT [23] falls into the category of decoder-only models. It undergoes pre-training on a Python corpus sourced from the CodeSearchNet dataset [27], comprising 1.1 million Python functions. For our purposes, we adopt the pre-trained CodeGPT-small version [^1] https://huggingface.co/microsoft/CodeGPT-small-py, which encompasses 124 million model parameters.
  • CodeT5: CodeT5 [4] belongs to the encoder-decoder model category and is similarly pre-trained on the Python corpus from the CodeSearchNet dataset. We select the pre-trained CodeT5-base version [^2] https://huggingface.co/Salesforce/codet5-base, which comprises 220 million model parameters.
  • CodeLlama: CodeLlama [12] represents another decoder-only model, specialized for code-related tasks and based on Llama2 [15]. It is pre-trained on an even larger Python corpus, encompassing a staggering 100 billion tokens sourced from a Python-centric dataset [12]. For our purposes, we adopt the pre-trained CodeLlama-7b version [^3] https://huggingface.co/codellama/CodeLlama-7b-Python-hf, featuring a substantial 7 billion model parameters.
We refer to the variants of ToolGen, namely ToolGen-gpt, ToolGen-t5, and ToolGen-llama, corresponding to the underlying base models CodeGPT, CodeT5, and CodeLlama, respectively.
Autocompletion Tool. We employ Jedi [19] as our autocompletion tool. Jedi is a static analysis tool designed for Python, commonly utilized within integrated development environments (IDEs) and editor plugins. Utilizing Jedi, ToolGen can trigger autocompletion, generating a list of suggestions that encompasses repository-level dependencies, including user-defined attributes and functions.
Trigger Insertion. To create the augmented dataset for fine-tuning the employed base model, we begin with the Python corpus from the CodeSearchNet dataset. Since the CodeSearchNet dataset does not provide complete code repositories from which to extract Python functions, we initiate the process by crawling the code repositories listed in the dataset. Subsequently, we follow the procedure outlined in Section III-B1 to extract and augment functions within these code repositories, ultimately generating the augmented dataset. It’s important to note that the CodeSearchNet dataset includes a partitioning into training, validation, and test sets. For our trigger insertion process, we exclusively utilize the code repositories associated with the training set. The resulting augmented dataset comprises a total of 256,870 pairs of docstrings and augmented functions, which are sourced from 12,231 distinct Python code repositories. Regarding dataset statistics, the average token count for docstrings is 10.96, and for augmented functions, it is 54.13. Additionally, the special token <COMP> appears an average of 5.37 times within these functions.
Model Fine-tuning. In the fine-tuning process, we adopt different strategies for CodeGPT, CodeT5, and CodeLlama: For CodeGPT and CodeT5, we perform full-parameter fine-tuning, optimizing all model parameters during this phase. In the case of CodeLlama, we employ LoRA with a reduction factor (r) of 16 and a scaling factor (alpha) of 16 to achieve parameter-efficient fine-tuning. This approach allows us to optimize only 0.25% of the trainable parameters in comparison to the original CodeLlama model. The fine-tuning settings for learning rate and batch size are consistent across all three models, with a learning rate of 5E-6 and a batch size of 32. However, the number of epochs differs: 10 epochs for CodeGPT and CodeT5, while CodeLlama undergoes fine-tuning for 2 epochs. To ensure reproducibility, we set the seed for random functions to 42 consistently across all packages and libraries used.
IV-C Evaluation Benchmark
IV-C1 Dataset
The construction of our benchmark dataset follows this process: We begin by crawling the code repositories listed in the test set of the CodeSearchNet dataset. Next, we extract pairs of docstrings and functions from these repositories by parsing and traversing Abstract Syntax Trees (ASTs), similar to the method described in Section III-B1. This process yields an evaluation dataset comprising 12,744 Python functions sourced from 680 code repositories. On average, the docstrings contain 10.63 tokens, while the functions consist of an average of 54.56 tokens.
IV-C2 Baselines
We develop three baseline approaches by fine-tuning these same base models but performing straightforward code generation without tool integration. The different variants of ToolGen and the corresponding baselines are presented in Table I, along with the base models they employ.
Specifically, the fine-tuning process for the baselines involves using the 256,870 pairs of docstrings and functions from the augmented dataset. Notably, the fine-tuning is conducted on the original functions, prior to the introduction of <COMP> tokens. The training configurations, including learning rates and training epochs, mirror those employed in the implementation of ToolGen. After fine-tuning, these models are utilized to perform straightforward code generation, as outlined in Section II-A. We label the three baseline approaches as follows:
  • Base-gpt: Represents straightforward code generation using the CodeGPT model fine-tuned on original functions.
  • Base-t5: Signifies straightforward code generation using the CodeT5 model fine-tuned on original functions.
  • Base-llama: Designates straightforward code generation with the CodeLlama model fine-tuned on original functions.
TABLE I: Variants of ToolGen and corresponding baselines.
  Approach Base Model
  Base-gpt CodeGPT
ToolGen-gpt (ours)
Base-t5 CodeT5
ToolGen-t5 (ours)
Base-llama CodeLlama
ToolGen-llama (ours)
IV-C3 Metrics
In our evaluation, we employ a combination of commonly used similarity metrics and introduce two novel metrics tailored specifically to repository-level code generation.
Similarity Metrics: We utilize the following well-established similarity metrics to measure the correspondence between generated functions and their ground-truth counterparts:
  • BLEU-4 [28]: This metric assesses the quality of generated code by comparing n-grams (sequences of n consecutive tokens) in the generated functions with those in the ground-truth functions.
  • CodeBLEU [23]: Specifically designed for code generation tasks, CodeBLEU evaluates the accuracy of code generation models by considering code-specific vocabulary and structure.
  • Edit Similarity (EditSim) [29]: This metric measures the similarity between two pieces of functions by analyzing the edit operations required to transform one into the other. It provides insights into code structure alignment.
Repository-level Metrics: To assess the effectiveness of both ToolGen and the baselines in repository-level code generation, we introduce the Repository-level metrics, namely Dependency Coverage (DepCov) and Error Rate (ErrRate).
  • Dependency Coverage (DepCov): This metric calculates the ratio of repository-level dependencies, including user-defined functions and attributes, that appear in ground-truth functions and are covered by the generated functions. The process involves static analysis using Jedi to identify expressions (such as function calls and attribute accesses) that can be traced back to their definitions within the code repository. We collect these expressions as a set and evaluate the extent to which they are addressed by the corresponding generated function. For example, if the attribute definition for “self._registered_updates” can be inferred for a ground-truth function, we examine whether it is properly addressed in the corresponding generated function.
  • Static Validity Rate (ValRate): As repository-level dependencies can potentially introduce dependency errors in generated functions, we introduce the Static Validity Rate metric (ValRate) to evaluate the effectiveness of ToolGen in reducing dependency errors. This metric evaluates the proportion of generated functions that successfully pass a static check for dependency errors, specifically no-member and undefined-variable. To perform this evaluation, we incorporate the generated functions into their respective code repositories and conduct static lint analysis using pylint [30]. Functions that do not exhibit syntax errors, no-member, or undefined-variable errors are deemed statically valid. It’s worth noting that this work does not address logic errors, as their identification typically necessitates the execution of corresponding unit tests.
V Results and Analyses
V-A RQ1: Similarity Metrics
The evaluation results of similarity metrics are presented in Table II. When comparing the performance of ToolGen’s variants and the three different base models, namely CodeGPT, CodeT5, and CodeLlama, we observe that ToolGen achieves similarity scores comparable to the baselines.
TABLE II: Evaluation results for similarity metrics. $\Delta$ indicates the metric improvement or reduction of ToolGen’s variants compared to the baselines.
  Approach Similairty Metrics
BLEU-4
CodeBLEU
EditSim
  Base-gpt
0.304
0.284
47.6%
ToolGen-gpt (ours)
0.315 ($\Delta=+3.6\%$)
0.275 ($\Delta=-3.2\%$)
47.1% ($\Delta=-1.1\%$)
Base-t5
0.355
0.291
63.7%
ToolGen-t5 (ours)
0.362 ($\Delta=+2.0\%$)
0.288 ($\Delta=-1.0\%$)
62.3% ($\Delta=-2.2\%$)
Base-llama
0.368
0.310
65.7%
ToolGen-llama (ours)
0.371 ($\Delta=+0.8\%$)
0.295 ($\Delta=-4.8\%$)
62.4% ($\Delta=-5.0\%$)
To provide a detailed breakdown, when utilizing CodeGPT as the base model, ToolGen-gpt demonstrates a 3.6% improvement in BLEU-4 compared to Base-gpt. However, it exhibits a 3.2% decrease in CodeBLEU and a 1.1% decrease in EditSim. Similarly, with the base model CodeT5, ToolGen-t5 exhibits a 2.0% enhancement in BLEU-4 relative to Base-t5 but experiences a 1.0% reduction in CodeBLEU and a 2.2% decrease in EditSim. In the case of the larger base model CodeLlama, ToolGen-llama shows an improvement of 0.8% in BLEU-4 compared to Base-t5 but encounters a 4.8% decrease in CodeBLEU and a 5.0% decrease in EditSim.
SUMMARY: ToolGen demonstrates competitive performance in similarity metrics with the baselines across various base models.
V-B RQ2.a: Dependency Coverage
Table III displays the evaluation results for repository-level Dependency Coverage (DepCov). Notably, our approach ToolGen demonstrates significant superiority over the baselines across all three base models. Specifically, when employing the base models CodeGPT, CodeT5, and CodeLlama, ToolGen surpasses the corresponding baselines in Dependency Coverage by 45.8%, 37.3%, and 15.2%, respectively.
These results underscore the effectiveness of the tool-integrated generation process in enhancing awareness of repository-level dependencies, a challenge often unaddressed by the conventional code LLMs. For instance, consider the incomplete function in Figure 1: in a straightforward CodeLlama generation, it fails to recognize the valid attributes of “self”. However, through tool-integrated generation, ToolGen leverages Jedi to deduce a list of completion suggestions, enabling it to select the most suitable one and cover target repository-level dependencies, including the usage of user-defined functions and attributes.
Despite the considerable improvement in repository-level Dependency Coverage facilitated by our approach, it is essential to acknowledge that the overall coverage remains limited. This limitation arises from the fact that code LLMs generate function tokens sequentially from left to right. Consequently, errors tend to accumulate as the token count increases due to the exposure bias problem [31, 32, 33]. This means that code LLMs often make incorrect token predictions at certain generation steps and may not produce <COMP> tokens to trigger autocompletion tools, especially for long functions.
SUMMARY: Our approach, ToolGen, consistently outperforms the baselines in repository-level Dependency Coverage across all three base models by ranging from 15.2% to 45.8%. These results highlight the effectiveness of our tool-integrated generation process in addressing the crucial issue of enhancing awareness of repository-level dependencies, which is often a challenge for conventional code LLMs in repository-level code generation.
TABLE III: Evaluation results for repository-level metrics. DepCov and ValRate represent Dependency Coverage and Static Validity Rate, respectively. $\Delta$ indicates the metric improvement or reduction of ToolGen’s variants compared to the baselines.
  Approach Repository-level Metrics
DepCov
ValRate
  Base-gpt
8.3%
32.2%
ToolGen-gpt (ours)
12.1% ($\Delta=+45.8\%$)
35.7% ($\Delta=+10.9\%$)
Base-t5
11.0%
48.6%
ToolGen-t5 (ours)
15.1% ($\Delta=+37.3\%$)
69.1% ($\Delta=+42.2\%$)
Base-llama
13.8%
47.8%
ToolGen-llama (ours)
15.9% ($\Delta=+15.2\%$)
67.1% ($\Delta=+40.4\%$)
V-C RQ2.b: Static Validity Rate
Table III also presents the evaluation results for Static Validity Rate (ValRate) in repository-level lint analysis, with a particular focus on no-member and undefined-variable errors. Remarkably, our approach, ToolGen, consistently exhibits significantly higher Static Validity Rate compared to the baselines across all three base models. Specifically, when employing the base models CodeGPT, CodeT5, and CodeLlama, ToolGen increases the Static Validity Rate by 10.9%, 42.2%, and 40.4%, respectively.
These results underscore the effectiveness of our tool-integrated generation process in mitigating the production of invalid identifiers during code generation within a specific repository context. For instance, let’s revisit the incomplete function in Figure 1: in a straightforward CodeLlama generation, it may predict a non-existent attribute, such as “updates”, for “self”. In contrast, through our tool-integrated approach, only valid completion suggestions inferred by Jedi are considered as candidates, thereby preventing numerous no-member and undefined-variable errors.
SUMMARY: Our approach, ToolGen, consistently achieves significantly higher Static Validity Rate in repository-level lint analysis compared to the baselines, with improvements ranging from 10.9% to 42.2%. These results underscore the effectiveness of our tool-integrated generation process in mitigating the generation of invalid identifiers, a common challenge faced by conventional code LLMs in the context of repository-level code generation.
V-D RQ3: Generalizability
Based on the results presented in Table II and Table III, our tool-integrated generation approach consistently improves performance in repository-level metrics while maintaining comparative similarity metrics across various model architectures (decoder-only and encoder-decoder) and parameter scales (ranging from 124 million to 7 billion).
Specifically, for the encoder-decoder model CodeT5, our tool-integrated generation approach demonstrates well-balanced and comprehensive enhancements in repository-level metrics. However, for the two decoder-only models, CodeGPT and CodeLlama, the improvements achieved by ToolGen vary between Dependency Coverage and Static Validity Rate. When employing the smaller CodeGPT as the base model, ToolGen significantly enhances Dependency Coverage but exhibits less pronounced improvements in the Static Validity Rate. In contrast, when employing the larger CodeLlama, the results are reversed. Upon closer examination of the generated functions, we attribute these variations to CodeGPT’s occasional difficulty in predicting the <EOS> token correctly, leading to syntax errors that violate the prerequisites for successful lint check. In the case of CodeLlama, the initially high Dependency Coverage, owing to its extensive parameter scale and training data, poses challenges for improvement through parameter-efficient fine-tuning (PEFT) with LoRA (only 0.25% trainable parameters).
SUMMARY: Our tool-integrated generation approach consistently improves repository-level metrics while maintaining competitive similarity metrics across various model architectures and parameter scales. This suggests that our approach is versatile and has the potential for broader applicability with other base models in repository-level code generation.
VI Discussion
VI-A Limitations
Static Autocompletion Tools for Dynamically Typed Programming Languages: Currently, our implementation of ToolGen is specific to Python, a dynamically typed programming language. However, the autocompletion tools used in ToolGen rely on static analysis, which can sometimes fail to trigger for certain repository-level dependencies. For instance, when the type of a function parameter cannot be explicitly inferred through static analysis, autocompletion tools may struggle to deduce attributes defined within the argument type. In the future, we plan to explore the integration of comprehensive type inference tools, such as learning-based tools, into the code generation process alongside autocompletion tools to enhance Python code generation.
Greedy Next Token Prediction in Generation Process: During the generation process, we employ a greedy strategy for next token prediction, where the token with the highest probability is selected using the argmax function. This greedy prediction strategy can occasionally lead the model to choose sub-optimal tokens for subsequent steps, resulting in code that may not be of the high quality. To address this issue, we intend to investigate the incorporation of techniques such as beam search and other advanced decoding methods into our tool-integrated generation process to mitigate the challenges posed by greedy prediction.
Repository-level Evaluation Metrics: In the computation of the two repository-level evaluation metrics, namely dependency coverage and success rate, we employ static analysis tools to identify target expressions and perform lint examinations. Similar to the autocompletion tools, these static tools may introduce a degree of inaccuracy into the calculated metrics. However, it is essential to note that this does not significantly impact the demonstrated effectiveness of ToolGen, as the baseline metrics are also determined using the same static tools. Another limitation is that our evaluation does not encompass metrics such as pass rate or pass@k for test cases, which are commonly used to assess the logical correctness of standalone functions generated by code LLMs. The primary reason for this omission is that such metrics typically rely on the execution of manually crafted unit test cases, which can be challenging to carry out in the context of repository-level code generation. Instead, our focus in this work centers on mitigating dependency errors, such as no-member and undefined-variable errors, stemming from the unawareness of repository-level dependencies.
VI-B Threats to Validity
Internal Threats: The first internal threat pertains to potential data quality issues common in learning-based approaches. To mitigate this threat, we construct our augmented dataset and evaluation benchmark dataset using the widely adopted CodeSearchNet dataset, which serves as a reliable source for pretraining and evaluating various code models. Another internal threat pertains to the potential data leakage for CodeLlama, as the code repositories in the benchmark dataset may have been encountered by CodeLlama during its pretraining phase. However, our generalizability evaluation (RQ3) provides evidence of consistent performance across the three ToolGen variants, suggesting that the improvements achieved by ToolGen-llama in repository-level code generation are not attributed to data leakage.
External Threats: Our implementation and evaluation of ToolGen are specific to the Python programming language. As a result, the findings may not be generalizable to other programming languages. Exploring the tool-integrated generation process for different languages is a valuable direction for future research.
VII Related Work
VII-A Code Generation
Code generation has long been a central focus in the field of software engineering. Recent developments have introduced a range of large language models for code (code LLMs), including Codex [3], CodeT5 [4], CodeT5+ [5], InCoder [6], AlphaCode [7], CodeGen [8], and CodeLlama [12], built upon the Transformer model architecture [14]. These models, either pretrained or fine-tuned on extensive code corpora, have the capability to automatically generate code based on provided natural language descriptions.
While these code LLMs have demonstrated significant effectiveness in generating standalone functions on existing benchmarks like HumanEval [3] and MBPP [16], they face substantial challenges when tasked with generating real-world functions within code repositories. The primary challenge stems from their lack of awareness of repository-level dependencies, such as user-defined functions and attributes, during the code generation process [17]. To address these challenges, researchers have proposed prompt engineering approaches to make code LLMs aware of repository-level dependencies. Shrivastava et al. [34] introduced the repository-level prompt generator, a framework for generating context-aware prompts without requiring access to the weights of the LLM. Bairi et al. [35] presented CodePlan, a task-agnostic framework that treats repository-level coding as a planning problem, using innovative techniques to generate multi-step code edits while considering context from the entire codebase, previous changes, and specific instructions.
In this study, we tackle the challenges associated with repository-level code generation by seamlessly integrating autocompletion tools into the generation process of code LLMs.
VII-B Incorporating External Tools into LLMs
Recent research [20, 21, 36, 37, 38, 39, 40, 41, 42, 43, 44] has explored the integration of external tools (e.g., search engines, web browsers, calculators, and python interpreters) into the LLM generation process, aiming to address their limitations in certain generation scenarios. For instance, Schick et al. propose ToolFormer [20], which augments datasets to instruct LLMs on invoking existing arithmetic calculators, effectively reducing errors in generated text related to arithmetic calculations. Building upon this idea, Zhang et al. introduce ToolCoder [21], designed to teach LLMs how to utilize information-retrieval-based (IR-based) API search tools during code generation. While ToolCoder is effective in generating functionally correct standalone functions, it falls short in addressing repository-level dependencies, limiting its ability to resolve dependency errors. Another relevant example is Repilot [22], which utilizes code completion tools to filter out impractical suggestions made by LLMs in the context of automatic program repairing (APR), focusing on generating valid bug-fix patches rather than entire functions.
In this paper, we integrate program-analysis-based autocompletion tools into the code LLM generation process to facilitate repository-level code generation.
VIII Conclusion
We present ToolGen, a novel approach that seamlessly integrates autocompletion tools into the code LLM generation process to effectively address repository-level dependencies. ToolGen encompasses two crucial phases: Data Augmentation and Model Fine-tuning, and Tool-integrated Code Generation. Our comprehensive evaluation showcases ToolGen’s substantial improvements in the two introduced repository-level metrics across three distinct code LLMs, while also demonstrating its competitiveness in widely-recognized similarity metrics. Moreover, our generalizability evaluation reaffirms ToolGen’s consistent performance when applied to diverse code LLMs, including various model architectures and scales.